Find the range of the function f defined by f(x y)=10-x^2-y^2 for all (x y) for which x^2+y^2<=9. Sketch two of its level curves.
1
Expert's answer
2020-05-06T18:50:07-0400

Function "f" is defined by "f(x, y)=10-x^2-y^2" for all "(x, y)" for which "x^2+y^2 \le 9." Also we know that for real numbers "x, y" we get "0 \le x^2+y^2."

Let us rewrite "f" in the form "f(x ,y)=10-(x^2+y^2)" .

Next we can write an inequation

"0\le x^2+y^2 \le 9."

Therefore,

"10-0 \ge 10-(x^2+y^2) \ge 10-9."

If we simplify the inequation above, we'll get

"10\ge 10-(x^2+y^2) \ge 1."

So "1 \le f(x,y) \le 10."


Let us discuss the properties of the range. We can see that points "(x,y)" are situated inside the circle with center "(0,0)" and radius "\sqrt{9}=3." This circle is shown by the black line in the figure below.

Now we'll consider the level curves. That means, we consider the set of points "(x,y)" where "f(x,y)" is a constant.


Let us consider the curve where "f(x,y) = 6" . We should find points "(x,y)" for which

"6=10-(x^2+y^2)" , so "x^2+y^2=4." The curve is also a circle with center in (0,0) and radius "\sqrt{4}=2." This circle is shown by the blue line in the figure below.


Let us consider the curve where "f(x,y)=9" . We should find points "(x,y)" for which "9=10-(x^2+y^2)", so "x^2+y^2=1." The curve is also a circle with center in (0,0) and radius "\sqrt{1}=1." This circle is shown by the red line in the figure below.